OpenStudy (yacoub1993):

Solve the equation for y. 5x – 4y = 12 A. y = -5/4x - 3 B. y = 4/5x + 3  C. y = 5/4x - 3 D. y = -4/5x + 3 

jimthompson5910 (jim_thompson5910):

The idea is to isolate y. So you need to subtract 5x from both sides. Then you need to divide both sides by -4 to completely isolate y 5x – 4y = 12 -4y = 12 - 5x -4y = -5x + 12 y = (-5x + 12)/(-4) y = (-5x)/(-4) + (12)/(-4) y = 5/4x - 3

OpenStudy (yacoub1993):

Graph the relation in the table. Then use the vertical-line test. Is the relation a function? x y –1 3 1 –1 2 2 1 3 A. The relation is a function. B. The relation is not a function. C. The relation is a function. D. The relation is not a function.